SPECTROSCOPY OF HIGHLY EXCITED HCNH^{+} USING INFRARED HETERODYNE VELOCITY MOD-ULATION

\underline{C. MICHAEL LINDSAY}, CHRISTOPHER F. NEESE, and TAKESHI OKA, Department of Chemistry, Department of Astronomy & Astrophysics, and the Enrico Fermi Institute, The University of Chicago, Chicago IL, 60637, USA.

Protonated hydrogen cyanide, HCNH^{+}, is very abundant in ionized environments containing hydrogen, carbon, and nitrogen. Motivated by its important role in interstellar chemistry as the precursor for HCN and HNC, Altman et al. ^{a} studied its fundamental C-H and N-H vibrational bands in 1984. Their analysis led to the detection of HCNH^{+} in interstellar space by Ziurys and Turner in 1985.^{b} Subsequent laboratory work concentrated on the measurement of the remaining fundamental bands, the pure rotational spectrum, and several isotopomers to determine the ground state molecular structure.

We have revisited this cation using an automated color center laser and a new highly sensitive technique that combines infrared heterodyne detection with velocity modulation,^{c} and recorded its high resolution spectrum in the C-H and N-H stretch regions (3000–3600 cm^{−1}). Highly excited HCNH^{+} was produced in a water cooled discharge of CH_{4}, N_{2}, H_{2}, and He. In this talk we report our analysis of this very dense spectrum and discuss the chemistry and dynamics of HCNH^{+} in our discharge with comparisons to HN_{2}^{+} and HCN.

aR. S. Altman, M. W. Crofton, and T. Oka, J. Chem. Phys, 80, 3911 (1984) and 81, 4255 (1984).

bL. M. Ziurys and B. E. Turner, Astrophys J., 302, L31 (1986).

cSee talk entitled “Infrared Laser Absorption Spectroscopy with Heterodyne Detection” .